OpenStudy (anonymous):

HELP ME PLEASE ! 1. x^2+y^2=64 c= r= 2. (x-1)^2+y^2=49\ c= r= 3. (x+3)^2+(y+9)^2=49 c= r= 4. x^2+y^2-4x-4y-6=0 c= r=

OpenStudy (gorv):

general form (x-a)^2 + (y-b)^2 = r^2 c=(a,b)

OpenStudy (gorv):

1.c=(0,0) r=8 2.c=(1,0) r=7 3.c=(-3,-9) r=7 4.make general form x^2-4x+4-4+y^2-4y+4-4-6=0 x^2-4x+4=(x-2)^2 y^2-4y+4=(y-2)^2 (x-2)^2 -4 +(y-2)^2 -4 -6=0 (x-2)^2+(y-2)^2-4-4-6=0 (x-2)^2+(y-2)^2=14 c=(2,2) r=14^(1/2)

OpenStudy (anonymous):

thank you @gorv !